OpenStudy (anonymous):

Assume a tractor has an initial price of $1000.00 and sells for $100.00 after 23 years. Assume that the tractor’s value depreciates linearly with the passing years. Let V represent the tractor's value in dollars and let t represent the number of years since the tractor was purchased. Write a formula to express the tractor's value as a function of time: V(t)=

OpenStudy (anonymous):

I know that V(t)= 1000-39.13t

OpenStudy (anonymous):

And I know that the vertical intercept of the value function V(t)= 1000 BUT I'm confused on how to get the horizontal intercept

OpenStudy (kropot72):

The horizontal intercept is found by plugging in zero for the value of V(t) and solving to find the value of t: 0 = 1000 - 39.13t ............(1) Now solve equation (1) to find the number of years for the tractor have zero value.

OpenStudy (anonymous):

t= -25.5558

OpenStudy (anonymous):

Now what?

OpenStudy (anonymous):

@zepdrix Can you help? :)

OpenStudy (kropot72):

There is an error in your solution. How did you get the negative sign?

OpenStudy (kropot72):

0 = 1000 - 39.13t ............(1) Add 39.13t to both sides. This gives: 39.13t = 1000 Now divide both sides by 39.13, to get the value of t in years. Then round to result.

OpenStudy (anonymous):

Ah, I see. Thanks! 25.56 (y)

OpenStudy (kropot72):

You're welcome :)